17. A computer with a 5 stage pipeline is measured an d has the following characteristies Instruction Type Avg. stall cycles/instructions % of instructions Branches Loads & Stores ALU ops Other what is the average CPI for the computer? What is the speedup for this machine? .4 .2 2 .2 .3 .5 a) b)
Answer #1

a) Avrage CPI= Σmaxtype ofinstructioni=1((Cycles taken per type of instruction )i * (Percentage of that instruction present)i );

Branch .3 .4
load .1 .2
ALU .5 .2
other .1 .2
  • Load (5 cycles)
  • Store (4 cycles)
  • R-type (4 cycles)
  • Branch (3 cycles)3.
  • Jump (3 cycles)

CPI=(0.3*3+0.1*5+0.5*4+0.1*1)=3.5

b)CPI from Avg. stall Cycles/instructions=0.4+0.2+0.2+0.2=1.0

Speedup for this machin=3.5 cpi / 1.0 cpi=3.5.